# Multi-Fidelity Bayesian Optimization With Across-Task Transferable Max-Value Entropy Search

### v2

@@ V Experiments
In this section, we empirically evaluate the performance of the proposed MFT-MES on the synthetic benchmark adopted in [29, 21], as well as on two real-world applications, namely radio resource management for wireless cellular systems [6, 7] and gas emission source term estimation [36].
V-A Benchmarks
@@ V-B Evaluation and Implementation
V-C Synthetic Optimization Tasks
For the first experiment, we consider synthetic optimization tasks defined by randomly generating Hartmann 6 functions [35]. Accordingly, the input domain is defined as $\mathcal{X}=[0,1]^{6}$, and the objective value $f^{(m)}_{n}(\mathbf{x})$ for a task $n$ at fidelity level $m$ is obtained as [21]
$f_{n}^{(m)}(\mathbf{x})=-\sum_{i=1}^{4}a_{i,m}\exp\bigg{(}-\sum_{j=1}^{6}\Delta_{i,j,n}A_{i,j}(x_{j}-P_{i,j})^{2}\bigg{)},$ (38)
where $a_{i,m}$, $A_{i,j}$ and $P_{i,j}$ are the $(i,m)$-th, $(i,j)$-th, and $(i,j)$-th entries, respectively, of matrices
$\mathbf{a}=\begin{pmatrix}1&1.01&1.02&1.03\\ 1.2&1.19&1.18&1.17\\ 3&2.9&2.8&2.7\\ 3.2&3.3&3.4&3.5\end{pmatrix},$
$\mathbf{A}=\begin{pmatrix}10&3&17&3.5&1.7&8\\ 0.05&10&17&0.1&8&14\\ 3&3.5&1.7&10&17&8\\ 17&8&0.05&10&0.1&14\end{pmatrix},$
$\text{and}\,\,\mathbf{P}=10^{-4}\begin{pmatrix}1312&1696&5569&124&8283&5886\\ 2329&4135&8307&3736&1004&9991\\ 2348&1451&3522&2883&3047&6650\\ 4047&8828&8732&5743&1091&381\end{pmatrix}.$ (39)
Optimization tasks differ due to the parameter $\Delta_{i,j,n}$ in (38) which are generated in an i.i.d. manner from the uniform distribution $\mathcal{U}(0.8,1.2)$. We set the cost levels as $\lambda^{(1)}=10,\lambda^{(2)}=15,\lambda^{(3)}=20,$ and $\lambda^{(4)}=25$, the total query cost budget in constraint (9) to $\Lambda=500$ for all tasks, and choose the observation noise variance in (6) as $\sigma^{2}=0.1$.
@@ V-C Synthetic Optimization Tasks
Fig. 3 shows the simple regret (40) as a function of the number of tasks $n$ observed so far. For MFT-MES, the weight parameter in (35) is set to $\beta=1.2$. Since MF-MES does not attempt to transfer knowledge across tasks, its average performance is constant for all values of $n$. By transferring knowledge across tasks, both Continual MF-MES and MFT-MES can reduce the simple regret as the number of observed tasks $n$ increases.
@@
V-D Radio Resource Management
@@
which evaluates the best function of the optimal value $f_{n}^{*}$ attained during the optimization process. Finally, we set $V=10$ particles for Continual MF-MES and MFT-MES. All other experimental settings are kept the same as in Sec. V-B.
In Fig. 5, we set the weight parameter $\beta=1.6$ for MFT-MES and plot the optimality ratio (41) as a function of the number of tasks, $n$. Confirming the discussions in Sec. V-C, the performance of MF-MES is limited by the lack of information transfer across tasks. In contrast, the performance of both Continual MF-MES and MFT-MES benefits from information transfer. Furthermore, MFT-MES outperforms Continual MF-MES after processing 12 tasks. Through a judicious choice of input and fidelity levels targeting the shared parameters $\bm{\theta}$, at the end of 40-th task, MFT-MES provides, approximately, an $7\%$ gain in terms of optimality ratio over Continual MF-MES, and a $23\%$ gain over MF-MES.
The impact of the weight parameter $\beta$ used by MFT-MES on the performance evaluated at the 20-th and 40-th task is illustrated in Fig. 6. MFT-MES with any weight parameter $\beta>0$ outperforms all other schemes given the same number of tasks observed so far. The best performance of MFT-MES is obtained for $\beta=1.2$ at $n=20$ tasks, and for $\beta=1.6$ at $n=40$ tasks. Therefore, as discussed in Sec. V-C, a larger value of weight parameter $\beta$ is preferable as the number of tasks $n$ increases. Moreover, the performance of MFT-MES is quite robust to the choice of $\beta$. For instance, selecting larger weights $\beta$, up to $\beta=2.4$, is seen to yield a mild performance degradation as compared to the best settings of weight parameters $\beta=1.2$ for $n=20$ and $\beta=1.6$ for $n=40$.
@@
V-E Gas Emission Source Term Estimation
Finally, we consider an application to the reverse problem formulation of gas emission source term estimation introduced in [36]. The problem aims to optimize a decision vector $\mathbf{x}$ that identifies the characteristics of the gas emission point source based on the Pasquill-Gifford dispersion model [55]. The feasible input domain of the vector is defined as $\mathbf{x}\in[10,5000]\times[-500,500]^{2}\times[0,10]\subset\mathbbm{R}^{4}$, with the first parameter being the source emission rate and the rest of the parameters describing the location of the emission source. Tasks are distinguished by the different locations of the sensors used to measure the concentration of emissions.
The objective function $f_{n}(\mathbf{x})$ is defined as the sum of the squared errors between the concentration measured at the sensors and the concentration calculated by the dispersion model given parameters $\mathbf{x}$. The fidelity of the evaluation of the objective function $f_{n}(\mathbf{x})$ depends on the atmospheric conditions, which can be classified into $M=6$ fidelity levels controlled by dispersion coefficients as in [36]. We set the cost levels as $\lambda^{(m)}=10+5\cdot(m-1)$; the total query cost budget is set to $\Lambda=750$ for every task; and the observation noise variance is set to $\sigma^{2}=10^{-3}$. The performance of all methods is measured by the simple regret in (40).
In Fig. 7, we set the weight parameter in (35) to $\beta=1.5$ for MFT-MES, and plot the simple regret (40) as a function of the number of tasks $n$ observed so far. The results demonstrate again the capacity of both Continual MF-MES and MFT-MES to transfer knowledge across tasks, achieving better performance as compared to MF-MES. After processing at least four tasks, MFT-MES outperforms Continual MF-MES. In particular, MFT-MES obtains a lower simple regret by a factor of two as compared to Continual MF-MES at the end of task $n=17$, confirming the importance of accounting for knowledge transfer in the acquisition function (35).
Figure 7: Gas emission source term estimation: Simple regret (40) against the number of tasks, $n$, for MF-MES, Continual MF-MES $(\beta=0)$, and MFT-MES ($\beta=1.5$) with $V=10$ particles.
In a manner similar to Sec. V-D, we demonstrate the impact of weight parameter $\beta$ on the simple regret evaluated at the 7-th and 17-th task in Fig. 8. The superiority of MFT-MES over all other schemes is observed to hold for any values of weight parameter $\beta>0$. MFT-MES achieves the best performance with weight parameter around $\beta=1.0$ for $n=7$, and approximately $\beta=1.5$ for $n=17$. The overall trend confirms the discussion in Sec. V-C, as a larger value of weight parameter $\beta$ is more desirable when the number of tasks $n$ increases.
VI Conclusion
In this work, we have have introduced MFT-MES, a novel information-theoretic acquisition function that balances the need to acquire information about the current task with goal of collecting information transferable to future tasks. The key mechanism underlying MFT-MES involves modeling transferable knowledge across tasks via shared inter-task latent variables, which are integrated into the acquisition function design and updated following Bayesian principles. From synthetic optimization tasks and real-world examples, we have demonstrated that the proposed MFT-MES scheme can obtain performance gains as large as an order of magnitude in terms of simple regret as compared to the state-of-art scheme that do not cater to the acquisition of transferable knowledge.
@@ VI Conclusion
Figure 8: Gas emission source term estimation: Simple regret (40) against weight parameter $\beta$, for MF-MES, Continual MF-MES with $n=7$ and $n=17$ tasks observed, and MFT-MES with $n=7$ and $n=17$ tasks observed. The number of particles for Continual MF-MES and MFT-MES is set to $V=10$.

### v3

@@ VI Experiments
In this section, we empirically evaluate the performance of the proposed MFT-MES on the synthetic benchmark adopted in [25, 18], as well as on a real-world application, namely radio resource management for wireless cellular systems [5, 6]. The Supplementary Material contains an additional experiment on a gas emission source term estimation problem [32].
VI-A Benchmarks
@@ VI-B Evaluation and Implementation
VI-C Synthetic Optimization Tasks
For the first experiment, we consider synthetic optimization tasks defined by randomly generating Hartmann 6 functions [31], the details of the experimental settings can be found in Supplementary Materials.
@@ VI-C Synthetic Optimization Tasks
Fig. 3 shows the simple regret (53) as a function of the number of tasks $n$ observed so far. For MFT-MES, the weight parameter in (37) is set to $\beta=1.2$. Since MF-MES does not attempt to transfer knowledge across tasks, its average performance is constant for all values of $n$. By transferring knowledge across tasks, both Continual MF-MES and MFT-MES can reduce the simple regret as the number of observed tasks $n$ increases, which confirms the theoretical result in Theorem 1.
@@
VI-D Radio Resource Management
@@
which evaluates the best function of the optimal value $f_{n}^{*}$ attained during the optimization process. Finally, we set $V=10$ particles for Continual MF-MES and MFT-MES. All other experimental settings are kept the same as in Sec. VI-B.
@@
In Fig. 5, we set the weight parameter $\beta=1.6$ for MFT-MES and plot the optimality ratio (54) as a function of the number of tasks, $n$. Confirming the discussions in Sec. VI-C and Theorem 1, the performance of MF-MES is limited by the lack of information transfer across tasks. In contrast, the performance of both Continual MF-MES and MFT-MES benefits from information transfer. Furthermore, MFT-MES outperforms Continual MF-MES after processing 12 tasks. Through a judicious choice of input and fidelity levels targeting the shared parameters $\bm{\theta}$, at the end of 40-th task, MFT-MES provides, approximately, an $7\%$ gain in terms of optimality ratio over Continual MF-MES, and a $23\%$ gain over MF-MES.
The impact of the weight parameter $\beta$ used by MFT-MES on the performance evaluated at the 20-th and 40-th task is illustrated in Fig. 6. MFT-MES with any weight parameter $\beta>0$ outperforms all other schemes given the same number of tasks observed so far. The best performance of MFT-MES is obtained for $\beta=1.2$ at $n=20$ tasks, and for $\beta=1.6$ at $n=40$ tasks. Therefore, as discussed in Sec. VI-C, a larger value of weight parameter $\beta$ is preferable as the number of tasks $n$ increases. Moreover, the performance of MFT-MES is quite robust to the choice of $\beta$. For instance, selecting larger weights $\beta$, up to $\beta=2.4$, is seen to yield a mild performance degradation as compared to the best settings of weight parameters $\beta=1.2$ for $n=20$ and $\beta=1.6$ for $n=40$.
VII Conclusion
In this work, we have have introduced MFT-MES, a novel information-theoretic acquisition function that balances the need to acquire information about the current task with goal of collecting information transferable to future tasks. The key mechanism underlying MFT-MES involves modeling transferable knowledge across tasks via shared GP model parameters, which are integrated into the acquisition function design and updated following Bayesian principles. From theoretical analysis and empirical results, we have demonstrated that the proposed MFT-MES can obtain performance gains as large as an order of magnitude in terms of simple regret as compared to the state-of-art scheme that do not cater to the acquisition of transferable knowledge.